# On Efficiency-Effectiveness Trade-off of Diffusion-based Recommenders

### v1

@@ 1 Introduction
Diffusion models [1, 2, 3] have recently shown strong potential in generative sequential recommendation [4, 5, 6, 4, 7], owing to their remarkable ability to model complex distributions of user behaviors and generate oracle items that best match user preferences. At the core of diffusion-based sequential recommendation [4, 5, 7] is framing a theoretical continuous process [2, 3, 8] which gradually reconstructs the oracle items from random noise. Typically, such theoretical continuous process (e.g., Stochastic Differential Equations (SDE) [2]) is realized by multi-step discrete approximations [9, 10] (e.g., Denoising Diffusion Probabilistic Models (DDPM) [1]) for practical implementation: they add noise to the ground-truth next items in the forward process and then denoise it step by step in the reverse process with denoising models, guided by the interaction history [11].
@@
While achieving success, this discrete approximation inherently introduces discretization errors [12, 13, 14] between the practical reverse trajectory (i.e., the dashed line of discrete DDPM in Figure 1(a)) and the theoretical trajectory (i.e., the solid line of continuous SDE in Figure 1(a)). Such discretization error is primarily caused by high-order truncation [13] in numerical solvers and accumulates across the multi-step reverse process. When approximating SDE [2] with finite steps $T$, discretization errors exhibit a growth as $T$ decreases [15, 16], which result in an efficiency-effectiveness trade-off: increasing the number of steps $T$ (typically 1,000 steps) for diffusion-based recommenders improves alignment between generated items and user preferences but incurs substantial inference overhead, while aggressively reducing total steps $T$ accelerates generation at the cost of amplified discretization error and misalignment. As shown in Figure 1(b), compared with traditional recommenders (e.g., SASRec [17] and PDRec [18]), the diffusion-based recommender, DreamRec [4] that requires 1,000 steps to achieve comparable performance, exhibits substantial inference overhead. Conversely, reducing $T$ significantly boosts acceleration yet leads to notable performance drops of DreamRec — evidenced by a significant NDCG drop from $T=2,000$ to $T=1$ on different datasets in Figure 1(c).
To resolve the trade-off, we propose TA-Rec, a two-stage framework that realizes one-step generation by smoothing the denoising function in pretraining stage and reduces the trajectory deviation by optimizing denoising models with user preferences in fine-tuning stage. Technically, in the pretraining stage, TA-Rec proposes temporal consistency regularization (TCR) to enforce consistency between denoising representations at adjacent timesteps (steps t1t-1 and $t$), smoothing the denoising function [19] to map noise to oracle items directly [16, 15] with bounded error. This could improve diffusion models’ efficiency in one-step generation without compromising the recommendation performance. In the fine-tuning stage, TA-Rec introduces adaptive preference alignment (APA) [7, 20] to mitigate the trajectory deviation and enhance recommenders’ effectiveness. Specifically, APA constructs preference pairs by selecting positive and negative items from the interaction space. Recognizing that the denoising trajectory from noise to oracle items inherently varies with both user preferences and noise magnitude, APA dynamically calibrates alignment strength according to both the similarity between preference pairs and the time steps of the noisy input, optimizing denoising models in a more refined manner. Notably, when the preference pair is hard to distinguish and the noisy input has a large noise degree, APA strategically reduces optimization strength to avoid overfitting to ambiguous preferences and random noise.
@@ 2 Realated Wrok
Accelerating Diffusion [15, 14] [15, 14] addresses the computational bottleneck inherent in the multi-step generation process of diffusion models. Recent advancements fall into two primary categories: training-free numerical solvers [15, 32] and training-based model distillation [33, 34, 35, 36]. The former optimizes numerical solvers to serve as faster samplers during the reverse process, reducing inference steps to 20-50 without tuning the original denoising model [15, 37, 38]. For instance, DPM++ [39] leverages adaptive high-order solvers to approximate the continuous SDEs [2], achieving faster generation via optimized numerical integration. To achieve more aggressive acceleration (1-4 steps), the latter approach focuses on tuning the denoising model through knowledge distillation. Specifically, consistency models [16, 40] distill pre-trained diffusion models into single-step generators by enforcing self-consistency of the ODE trajectory. However, these methods may either rely on well-pretrained models or a deterministic ODE process, limiting their applicability to recommendation tasks due to the lack of unified recommendation benchmarks and the dynamic nature of user preferences. Instead, our work achieves one-step generation of diffusion-based recommenders by designing a temporal consistency loss for the stochastic DDPM process.
Aligning Diffusion Models [20, 41, 42, 43] aim to optimize diffusion models using human preference data, an area that remains relatively underdeveloped compared to preference optimization in large language models (LLMs) [44, 45]. Current approaches often adapt existing LLM preference optimization algorithms to diffusion models. For example, Diffusion-DPO [20] takes the DPO loss [44] from LLMs and adapts it on diffusion models directly to improve image generation based on user preference. DSPO [43] aligns diffusion models with human preferences by distilling the score function of preferred image distributions into the model’s pretrained score functions, leveraging score matching. Furthermore, Diffusion-NPO [42] approaches preference alignment by training an additional model to specifically model negative preferences. To enable alignment at a more fine-grained level, our work designs an adaptive preference optimization algorithm that aligns the generation of diffusion-based recommenders according to both timesteps and pair similarity.
@@ 3.1 Diffusion-based Sequential Recommendation
where $\mathbf{g}$ is the guidance signal extracted from users’ historical interaction sequences $\mathbf{x}^{1:N-1}$ with a Transformer [46] following [4, 7], $f_{\theta}(\cdot)$ is a denoising model parameterized by MLP, reconstructing the target item $\mathbf{x}$ as $\mathbf{\hat{x}}_{0}$ directly under the guidance $\mathbf{g}$, as shown in the green part of Figure 2. Commonly, classifier-free guidance paradigm [11] can be utilized by replacing the guidance $\mathbf{g}$ with a dummy token $\Phi$ at probability $\rho$, allowing for the integration of conditional and unconditional training. During inference, the pure Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$ serves as the input, then the denoising model $f_{\theta}(\cdot)$ denoises items as $\mathbf{\hat{x}_{0}}$ under the guidance $\mathbf{g}$. The iterative reverse process for generation is:
@@ 3.2 Direct Preference Optimization on Diffusion Models
Direct Preference Optimization (DPO) [44] is a reward-model-free method that aligns LLMs with human preferences via a supervised loss derived from pairwise comparison. To adapt DPO [44] to the diffusion model [20, 43, 42], the key idea is to increase denoising models’ likelihood of preferred generation while decreasing that of dispreferred ones. We assume the pairwise preference data as $D={(\mathbf{x}^{+},\mathbf{x}^{-})}$, where $\mathbf{x}^{+}$ is preferred over $\mathbf{x}^{-}$. Following [20], the DPO training loss on diffusion model can be formulated as:
@@ 7 Acknowledgement
This research is supported by the National Natural Science Foundation of China (62572449, 62302321). This research is also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
@@ E.2 Broader Impact
E.3 Limitations
Training Cost: While TA-Rec achieves significant inference efficiency through one-step generation, its two-stage training framework (pretraining with TCR and fine-tuning with APA) requires more computational resources compared to end-to-end approaches. Additionally, the TCR necessitates executing the denoising model twice for the adjacent steps per training iteration, thereby incurring additional computational overhead during the pretraining stage.
Negative Sampling Strategy: The APA in the fine-tuning stage relies on in-batch random negative sampling, which may limit its ability to select hard negative items and catch user preferences. Integrating advanced sampling approaches (e.g., hardness-aware or dynamic adversarial sampling) could refine preference alignment during optimization.

### v3

@@ 1 Introduction
Diffusion models [1, 2, 3] have recently shown strong potential in generative sequential recommendation [4, 5, 6, 4, 7], owing to their remarkable ability to model complex distributions of user behaviors and generate oracle items that best match user preferences. At the core of diffusion-based sequential recommendation [4, 5, 7] is framing a theoretical continuous process [2, 3] which gradually reconstructs the oracle items from random noise. Typically, such theoretical continuous process (e.g., Stochastic Differential Equations (SDE) [2]) is realized by multi-step discrete approximations [8, 9, 10, 11] (e.g., Denoising Diffusion Probabilistic Models (DDPM) [1]) for practical implementation: they add noise to the ground-truth next items in the forward process and then denoise it step by step in the reverse process with denoising models, guided by the interaction history [12].
@@
While achieving success, this discrete approximation inherently introduces discretization errors [13] between the practical reverse trajectory (i.e., the dashed line of discrete DDPM in Figure 1(a)) and the theoretical trajectory (i.e., the solid line of continuous SDE in Figure 1(a)). Such discretization error is primarily caused by high-order truncation in numerical solvers and accumulates across the multi-step reverse process. When approximating SDE [2] with finite steps $T$, discretization errors grow as $T$ decreases [14, 15], creating an efficiency-effectiveness trade-off: increasing the number of steps $T$ (typically 1,000 steps) for diffusion-based recommenders improves alignment between generated items and user preferences but incurs substantial inference overhead, while aggressively reducing total steps $T$ accelerates generation at the cost of amplified discretization error and misalignment. As shown in Figure 1(b), compared with traditional recommenders (e.g., SASRec [16] and PDRec [17]), the diffusion-based DreamRec [4] requires 1,000 steps to achieve comparable performance, significantly slowing inference. Conversely, reducing $T$ significantly boosts acceleration yet leads to notable performance drops of DreamRec — evidenced by a significant NDCG drop from $T=2,000$ to $T=1$ on different datasets in Figure 1(c).
To resolve the trade-off, we propose TA-Rec, a two-stage framework that realizes one-step generation by smoothing the denoising function in pretraining stage and reduces the trajectory deviation by optimizing denoising models with user preferences in fine-tuning stage. Technically, in the pretraining stage, TA-Rec proposes temporal consistency regularization (TCR) to enforce consistency between denoising representations at adjacent timesteps, smoothing the denoising function [18] to map noise to oracle items directly [15, 14] with bounded error. This could improve diffusion models’ efficiency in one-step generation without compromising the recommendation performance. In the fine-tuning stage, TA-Rec introduces adaptive preference alignment (APA) [7, 19] to mitigate the trajectory deviation and enhance recommenders’ effectiveness. Specifically, APA constructs preference pairs by selecting positive and negative items from the interaction space. Recognizing that the denoising trajectory from noise to oracle items inherently varies with both user preferences and noise magnitude, APA dynamically calibrates alignment strength according to both the similarity between preference pairs and the time steps of the noisy input, optimizing denoising models in a more refined manner. Notably, when the preference pair is hard to distinguish and the noisy input has a large noise degree, APA strategically reduces optimization strength to avoid overfitting to ambiguous preferences and random noise.
@@ 2 Realated Wrok
Accelerating Diffusion [14, 31, 32, 33, 34] addresses the computational bottleneck inherent in the multi-step generation process of diffusion models. Recent advancements fall into two primary categories: training-free numerical solvers [14, 35, 36] and training-based model distillation [37, 38, 39, 40]. The former optimizes numerical solvers to serve as faster samplers during the reverse process, reducing inference steps to 20-50 without tuning the original denoising model [14, 41, 42]. For instance, DPM++ [43] leverages adaptive high-order solvers to approximate the continuous SDEs [2], achieving faster generation via optimized numerical integration. To achieve more aggressive acceleration (1-4 steps), the latter approach focuses on tuning the denoising model through knowledge distillation. Specifically, consistency models [15, 44] distill pre-trained diffusion models into single-step generators by enforcing self-consistency of the ODE trajectory. However, these methods may either rely on well-pretrained models or a deterministic ODE process, limiting their applicability to recommendation tasks due to the lack of unified recommendation benchmarks and the dynamic nature of user preferences. Instead, our work achieves one-step generation of diffusion-based recommenders by designing a temporal consistency loss for the stochastic DDPM process.
Aligning Diffusion Models [19, 45, 46, 47] aim to optimize diffusion models using human preference data, an area that remains relatively underdeveloped compared to preference optimization in large language models (LLMs). Current approaches often adapt existing LLM preference optimization algorithms to diffusion models. For example, Diffusion-DPO [19] takes the DPO loss [48] from LLMs and adapts it on diffusion models directly to improve image generation based on user preference. DSPO [47] aligns diffusion models with human preferences by distilling the score function of preferred image distributions into the model’s pretrained score functions, leveraging score matching. Furthermore, Diffusion-NPO [46] approaches preference alignment by training an additional model to specifically model negative preferences. To enable alignment at a more fine-grained level, our work designs an adaptive preference optimization algorithm that aligns the generation of diffusion-based recommenders according to both timesteps and pair similarity.
@@ 3.1 Diffusion-based Sequential Recommendation
where $\mathbf{g}$ is the guidance signal extracted from users’ historical interaction sequences $\mathbf{x}^{1:N-1}$ with a Transformer following [4, 7], $f_{\theta}(\cdot)$ is a denoising model parameterized by MLP, reconstructing the target item $\mathbf{x}$ as $\mathbf{\hat{x}}_{0}$ directly under the guidance $\mathbf{g}$, as shown in the green part of Figure 2. Commonly, classifier-free guidance paradigm [12] can be utilized by replacing the guidance $\mathbf{g}$ with a dummy token $\Phi$ at probability $\rho$, allowing for the integration of conditional and unconditional training. During inference, the pure Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$ serves as the input, then the denoising model $f_{\theta}(\cdot)$ denoises items as $\mathbf{\hat{x}_{0}}$ under the guidance $\mathbf{g}$. The iterative reverse process for generation is:
@@ 3.2 Direct Preference Optimization on Diffusion Models
Direct Preference Optimization (DPO) [49, 50, 51] is a reward-model-free method that aligns LLMs with human preferences via a supervised loss derived from pairwise comparison. To adapt DPO [48] to the diffusion model [19, 47, 46], the key idea is to increase denoising models’ likelihood of preferred generation while decreasing that of dispreferred ones. We assume the pairwise preference data as $D={(\mathbf{x}^{+},\mathbf{x}^{-})}$, where $\mathbf{x}^{+}$ is preferred over $\mathbf{x}^{-}$. Following [19], the DPO training loss on diffusion model can be formulated as:
@@ 7 Acknowledgement
This research is supported by the National Natural Science Foundation of China (62572449, 62302321). This research is also supported by the Fundamental Research Funds for the Central Universities (WK2100250065) and the advanced computing resources provided by the Supercomputing Center of the USTC.
@@ E.2 Broader Impact
E.3 Limitations and Future Work
Training Cost Limitation: While TA-Rec achieves significant inference efficiency through one-step generation, its two-stage training framework (pretraining with TCR and fine-tuning with APA) requires more computational resources compared to end-to-end approaches. Additionally, the TCR necessitates executing the denoising model twice for the adjacent steps per training iteration, thereby incurring additional computational overhead during the pretraining stage.
Future Work: We will explore integrating large language model (LLM) scaling laws with diffusion-based generative recommendation frameworks to develop a diffusion LLM for generative recommendation. Specifically, we aim to investigate how LLM emergent capabilities—such as context understanding and semantic reasoning—scale with model size, data volume, and computational resources to enhance the expressive power of diffusion processes in capturing complex user-item interaction patterns.